OpenStudy (anonymous):

Find the vertex of the quadratic equation: y = 2x2 + 12x − 4.

OpenStudy (anonymous):

First identify a, b, and c terms of the equations. a = 2, b = 12, c = -4

OpenStudy (anonymous):

To find x-coordinate of the vertex use the formula x = -b/2a

OpenStudy (anonymous):

To find y-coordinate of the vertex use this formula: (c-b^2)/4a

OpenStudy (anonymous):

i dont understand im sorry

OpenStudy (anonymous):

Do you get the a, b, and c part, just identifying them?

OpenStudy (anonymous):

yeah kinda

OpenStudy (anonymous):

a is just the number in front of x^2. b the number in front of x and c the number with no x. You cool with that?

OpenStudy (anonymous):

yeahh

OpenStudy (anonymous):

Okay so, do you know what a vertex is?

OpenStudy (anonymous):

yeahh

OpenStudy (anonymous):

Okay so you need to find the x and y coordinate of the vertex. and it's given by those two formulas that I gave you. So you can just plug the number we have for a, b, and c into those formulas, yeah?

OpenStudy (anonymous):

so x = -(12)/2(2) , yeah?

OpenStudy (anonymous):

yess

OpenStudy (anonymous):

Cool, so you can find the y-coordinate using the second formula, yes? :) what do you get?

OpenStudy (anonymous):

-3 -22 ?

OpenStudy (anonymous):

I got -3 for the x coordinate by my answer for y-coordinate is different. Remember c is -4 not positive 4.

OpenStudy (anonymous):

so y coordinate = (-4-12^2)/(4(2))

OpenStudy (anonymous):

-3 - 50 ?

OpenStudy (anonymous):

12^2 = 144

OpenStudy (anonymous):

-4 - 144 = -148

OpenStudy (anonymous):

4*2 = 8

OpenStudy (anonymous):

remember order of operations. Do what's in parenthesis first.

OpenStudy (anonymous):

-3 -58 ?

OpenStudy (anonymous):

y coordinate is = -148/8

OpenStudy (anonymous):

the formula is (c-b^2)/4a first you calculate what's inside of the parenthesis. then you calculate the bottom: 4a. Then you do the top divided by the bottom. (4a).

OpenStudy (anonymous):

so what's -148/8?